### v1

# Tokenizer Choice For LLM Training: Negligible or Crucial?

Abstract
The recent success of Large Language Models (LLMs) has been predominantly driven by curating the training dataset composition, scaling of model architectures and dataset sizes and advancements in pretraining objectives, leaving tokenizer influence as a blind spot. Shedding light on this underexplored area, we conduct a comprehensive study on the influence of tokenizer choice on LLM downstream performance by training 24 mono- and multilingual LLMs at a 2.6 B parameter scale, ablating different tokenizer algorithms and parameterizations. Our studies highlight that the tokenizer choice can significantly impact the model’s downstream performance, training and inference costs. In particular, we find that the common tokenizer evaluation metrics fertility and parity are not always predictive of model downstream performance, rendering these metrics a questionable choice for tokenizer evaluation. Furthermore, we show that multilingual tokenizers trained on the five most frequent European languages require vocabulary size increases of factor three in comparison to English. While English-only tokenizers have been applied to the training of multi-lingual LLMs in the past, we find that this approach results in a severe downstream performance degradation and additional training costs of up to 68%, due to an inefficient tokenization vocabulary.
†Equal contribution.
1 Introduction
LLMs have shown impressive capabilities in many downstream tasks in a zero/few-shot setting such as summarization, reading comprehension, translation, and commonsense reasoning Brown et al. (2020b); Touvron et al. (2023). To train a LLM, the currently established approach is to employ a tokenizer that splits the training documents into tokens where a token represents a word Bengio et al. (2000), a sub-word Schuster and Nakajima (2012); Sennrich et al. (2015); Wang et al. (2020), or a single character Gao et al. (2020b), and each token is represented in the model by an embedding vector that can be further processed.
The quality of a tokenizer can be assessed intrinsically and extrinsically. An intrinsic evaluation solely addresses the characteristics of tokenizers and their generated output in isolation, whereas the extrinsic evaluation measures the impact of the tokenizer on a downstream component, e.g., the Large Language Model (LLM) model.
While many different tokenization approaches have been proposed, ranging from character-based to word-based methods, the potential impact of different tokenizers is neglected w.r.t. LLMs, especially in the context of multilingual LLMs. Recent work proposed by Petrov et al. (2023) demonstrates that carelessly designed tokenizers applied to the training of multilingual LLMs result in severe inequalities and limitations across languages. In fact, text passages translated into different languages resulted in tokenized sequences that differ in length up to a factor of 15, affecting inference costs and latency during inference. Furthermore, it is known that an essential property for effectively learning transformer-based LLMs is the learning of long-range dependencies Vaswani et al. (2017). Given a fixed sequence length, learning to relate words far apart in the input text is impossible for languages whose text is excessively fragmented by the tokenizer.
Despite the importance of tokenizers and the potentially severe impact of poorly performing tokenizers, there exists so far no extensive study that holistically investigates the intrinsic and extrinsic tokenizer performance in a monolingual and multilingual setting with a focus on decoder-only models, which represent the backbone of current LLMs.
Therefore, we address this gap and conduct an extensive study in which we measure the impact of the tokenizer on the model performance. In particular, we make the following contributions:
  • We conduct a study investigating the intrinsic tokenizer performance.
  • We conduct a study investigating the extrinsic tokenizer performance, i.e., the impact of the tokenizer on the model’s downstream performance.
  • Investigate whether a correlation between the intrinsic and the extrinsic tokenizer performance exists.
2 Related Work
This section provides an overview of tokenization algorithms and their usage in encoder- and decoder-only transformer models.
2.1 Tokenization Approaches
Word Tokenization.
The most basic tokenization approach is to split sequences based on white spaces and consider each word as a token Bengio et al. (2000).
Subword tokenization.
Subword tokenization algorithms are data-driven tokenization approaches which can decompose words into subwords/multiple tokens and currently represent the established tokenization approach upon which LLMs rely Kudo and Richardson (2018); Petrov et al. (2023). Because subword tokenizers decompose words into subwords, they can process out-of-vocabulary words by merging subwords from the vocabulary Kudo and Richardson (2018). Examples of popular subword tokenizers are WordPiece Schuster and Nakajima (2012), BPE Gage (1994); Sennrich et al. (2015), Byte-Level BPE (BBPE) Wang et al. (2020), and Unigram Kudo (2018).
Character Tokenization.
Tokenization can also be performed on a character level or based on UTF-8 bytes. However, this results in an increased sequence length, which becomes computationally expensive in the transformer architecture, the current predominated architecture for LLMs due to the quadratic complexity of the self-attention layer in the sequence length Vaswani et al. (2017). Though, several approaches have been proposed to address this limitation Gao et al. (2020b); Tay et al. (2021); Xue et al. (2022); Clark et al. (2022); Yu et al. (2023).
2.2 Tokenizers in Transformers Models
Tokenizers in Encoder Models
Most research on tokenization has been conducted on encoder models. Rust et al. (2021) investigated whether the tokenizer choice impacts the downstream performance of multi- and monolingual BERT Devlin et al. (2018) models. Zhang et al. (2022) showed that better machine translation performance is often obtained when languages are equally sampled during the tokenizer training. Toraman et al. (2023) trained several medium-sized language models for Turkish and suggested that different subword tokenizers perform roughly equivalent, whereas word- and character-level tokenizers perform drastically worse on downstream tasks. Finally, Chirkova and Troshin (2022) analyzed the effect of employing different tokenizations on code-related tasks and demonstrated that carefully configured tokenizers could reduce average sequence length up to 40% or allow for small downstream performance improvements by up to 2% at a lower compression rate.
Tokenizers in Decoder Models
An overview of current mono- and multilingual LLMs is provided in Lin et al. (2022); Shliazhko et al. (2022); Scao et al. (2022). As part of their work, Shliazhko et al. (2022) ablated different tokenizer pre-processing approaches while keeping the tokenizer algorithm, the vocabulary size and the employed implementation fixed. In none of the other major LLMs works the extrinsic tokenizer performance has been studied.
3 Approach
To investigate the tokenizer impact on the model performance, we conducted an extensive ablation study. In detail, we created dedicated datasets for the training of the tokenizers and the models, trained BPE and Unigram tokenizers, and for each tokenizer we trained decoder-only models with a size of 2.6B parameters while keeping the remaining configuration (i.e., dataset and model hyper-parameters) fixed. This allowed us to measure the tokenizer’s impact on the model’s downstream performance in isolation.
3.1 Data
While creating our tokenizer and model training datasets, we ensure that the mixture proportions of data domains (Wikipedia, books, web text) follow the same distribution to avoid a domain shift between tokenizers training and model training. We created two datasets with 70B words where one of the datasets is monolingual, containing English documents, and the second is a multilingual dataset comprised of English, German, French, Italian, and Spanish documents. Our datasets are filtered and deduplicated and consist of web-crawled data (80%) and curated data (20%), comparable to related datasets used to train LLMs. In the multilingual dataset, the amount of web-crawled data is equally distributed across languages in terms of number of words. Further details about our data pipeline and the data composition are described in Table 5.
3.2 Tokenizer
Our studies rely on the two established tokenization algorithms, BPE and Unigram, and their implementation in the Huggingface tokenizer library Moi and Patry (2023) and the SentencePiece library Kudo and Richardson (2018). We considered both libraries in order to investigate the effect of differences in the pre-and post-processing steps and potential differences in the implementations. Due to missing pre-processing options for Huggingface’s Unigram implementation, which causes a large discrepancy in the resulting vocabulary compared to SentencePiece’s implementation of Unigram, we omitted the training of Unigram tokenizers based on Huggingface. Overall, we trained 24 different tokenizers, where one-half of the tokenizers were monolingual English tokenizers, and the other half of the tokenizers were multilingual tokenizers. Besides the tokenizer algorithm, language composition, and employed tokenizer library, we also varied the vocabulary size. Concrete tokenizer configurations are described in the Appendix B.
3.3 Models
To measure the impact of our trained tokenizers on the model downstream performance, we trained one model for each tokenizer. In particular, for each of our 24 trained tokenizers, we trained a 2.6B transformer-based decoder-only model on up to 52B tokens following the scaling law proposed by Hoffmann et al. (2022a). Additionally, serving as baselines, we trained a monolingual and a multilingual model using the pre-trained GPT-2 tokenizer. All models have been trained based on the causal language modeling training objective.
3.4 Evaluation
To assess the impact of the tokenizers on the model downstream performance, we first performed an intrinsic tokenizer evaluation, followed by an extrinsic evaluation, and finally, we investigated whether a correlation between both evaluation approaches is given.
The intrinsic evaluation aims to assess the generated output of tokenizers based on fertility and parity. Furthermore, the tokenizer’s vocabulary overlap with other tokenizers is computed. The intrinsic evaluation does not assess the impact of tokenizers on the model performance.
Fertility, the most common metric to evaluate a tokenizer’s performance Scao et al. (2022); Stollenwerk (2023); Rust et al. (2021), is defined as the average number of tokens that are required to represent a word or document. For a tokenizer $T$ and dataset $A$, the fertility can be calculated as the number of tokens in $A$ (when $T$ is applied) divided by the number of words in $A$. We calculate the fertility on a held-out set (10,000 documents), which was not used for the tokenizer training. For calculating the words of a document, we used whitespace splitting. Higher fertility scores correspond to weaker compression capabilities of the tokenizer.
Parity Petrov et al. (2023), which has been recently proposed, assesses how fairly a tokenizer treats equivalent sentences in different languages. A tokenizer $T$ achieves parity for language $A$ with respect to language $B$ if $\frac{|T(s_{A})|}{|T(s_{B})|}\approx 1$, where $s_{A}$ and $s_{B}$ denote the sets of all sentences in the corpora of language $A$ and $B$, respectively. We use the FLORES-200 Goyal et al. (2022) parallel corpus, consisting of the same sentences human-translated into 200 languages. We calculate the parity values for each tokenizer and the four non-English languages with respect to English (see Fig. 2 for an overview).
The extrinsic evaluation aims to explicitly assess the impact of a tokenizer on the model’s downstream performance. We selected a comprehensive set of downstream tasks covering commonsense reasoning, reading comprehension, translation, and math tasks to measure the downstream performance. For an exhaustive lists of all tasks, we refer the reader to Table 11 in Appendix E.
Additionally, we computed the impact of a tokenizer on the average computational costs of a given model per word during training and inference. The computational costs during training for one step including the forward and the backward pass can be estimated by
$C=96Bslh^{2}\left(1+\dfrac{s}{6h}+\dfrac{V}{16lh}\right),$ (1)
given a model with batch size $B$, sequence length $s$, $l$ layers, hidden size $h$ and vocabulary size $V$ Narayanan et al. (2021). The costs per token can be derived by $C_{\text{token}}=C/Bs$ and the average costs per word by $C_{\text{word}}=C_{\text{token}}\cdot\text{fertility}$.
4 Intrinsic Tokenizer Evaluation
In our intrinsic evaluation, we first compare the fertility and parity of the trained tokenizers (Section 4.1) and the overlap of their vocabularies (Section 4.2).
4.1 Fertility & Parity
(a) Non-English, multilingual documents
(b) English documents
Figure 1: Comparison of fertility scores between monolingual (English) tokenizer and multilingual tokenizers applied to (a) English and (b) non-English, multi-lingual documents.
Applying the described fertility and parity evaluation to the mono-/multilingual tokenizers, our analysis highlights the following two major aspects, as visible in Fig. 1 and Fig. 2.
Firstly, it can be observed that applying a monolingual tokenizer to multilingual data results in significantly higher fertility and parity scores (see Fig. 0(a) and Fig. 2). While multilingual tokenizers have lower fertility than monolingual English tokenizers on all non-English documents by a large margin, they are only slightly worse on tokenizing English documents, as shown in Fig. 0(b).
Secondly, with increasing vocabulary size, fertility and parity reduce in all cases, which can be explained by the tokenizer requiring fewer sub-word tokens when tokenizing text given a larger vocabulary. However, it can be observed that for monolingual English tokenizers, the fertility is less dependent on the vocabulary when tokenizing English documents, implying that 33k might be a sufficiently large vocabulary. This is especially interesting considering the fact that the LLaMA/LLaMA2 models have been trained with a vocabulary size of 33k Touvron et al. ; Touvron et al. (2023).
Figure 2: Comparison of parity scores between monolingual (English) tokenizer and multilingual tokenizers applied multi-lingual documents.
4.2 Vocabulary Overlap
To analyze the tokenizer similarity, we calculated the vocabulary overlap. Particularly, we assess Huggingface’s and SentencePiece’s BPE implementations. Their vocabulary overlap is depicted in Table 1.
The overlap is roughly constant across different vocabulary sizes, and the total overlap tends to be rather low, despite being the identical algorithm only implemented by two different libraries. Consequently, the tokenizers produce different tokenized sequences, possibly affecting model training and downstream performance. Investigating the underlying reasons, the low overlap might be attributed to different configuration and pre-processing options in these libraries. Due to the larger thesaurus in multilingual documents, the overlap for the multilingual tokenizer is lower than for the English tokenizers.
33k 50k 82k 100k
English 0.77 0.76 0.74 0.74
Multilingual 0.62 0.62 0.62 0.61
Table 1: Vocabulary overlap between the HuggingFace and SentencePiece BPE tokenizer for different vocab sizes.
(a) Non-English documents
(b) English documents
(c) German documents
Figure 3: Average compute (GFLOPs) required to process a single word within (a) multilingual, (b) English, and (c) German documents within a full training pass (including the backward pass).
5 Extrinsic Tokenizer Evaluation
In the following, we describe the results of our extrinsic evaluation of tokenizers. Section 5.1 describes the experimental setup, Section 5.2 presents the downstream performance of the trained models based on the investigated tokenizers, and Section 5.3 analyses the computational costs associated with each tokenizer when employed in a specific model.
5.1 Experimental Setup
To assess the impact of the tokenizers on the model downstream performance, we trained a decoder-only transformer model of size 2.6 B for each tokenizer. We trained our models for 52.6 B tokens following the scaling law proposed by Hoffman et al. Hoffmann et al. (2022b) based on the causal language modeling training objective. The hyper-parameters are described in Table 8 in the Appendix C.
We evaluated our models in zero-shot settings on a wide range of tasks, covering commonsense reasoning (HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2020), ARC easy and challenge Clark et al. (2018)), world knowledge (TriviaQA Joshi et al. (2017)), reading comprehension (LAMBADA Paperno et al. (2016) and their machine translated versions for German, Spanish, French, and Italian, RACE Lai et al. (2017), BoolQ Clark et al. (2019)), math (MATH Hendrycks et al. (2021)), multilingual natural language inference (XNLI Conneau et al. (2018)).
5.2 Downstream Performance
We split our analysis of the downstream performance into several parts. First, we discuss the overall results obtained for the investigated tokenizers, followed by presenting the impact of the tokenizer library (Section 5.2.1), the impact of the tokenizer algorithm (Section 5.2.2), and the impact of the vocabulary size (Section  5.2.3).
(a) Monolingual tokenizers.
(b) Multilingual tokenizers.
Figure 4: Top-1/-2/-3 count for mono-/multilingual tokenizers across mono-/multilingual tasks the tasks were evaluated on the accuracy metric.
Model DE FR IT ES EN
EN BPE-HF - - - - 45.44
BPE-SP - - - - 46.31
UNI-SP - - - - 46.01
MULTI BPE-HF 36.72 36.55 35.34 42.28 44.45
BPE-SP 36.96 36.95 36.12 41.96 44.92
UNI-SP 37.38 37.05 35.70 42.56 44.21
Table 2: Impact of the tokenizer algorithm and the tokenizer training library, scores are averaged over the vocabulary sizes.
Vocab DE FR IT ES EN
EN 32 - - - - 46.34
50 - - - - 46.30
82 - - - - 45.29
100 - - - - 45.88
MULTI 32 36.72 36.22 35.20 41.76 44.71
50 36.57 36.34 34.60 42.22 43.91
82 37.40 37.22 36.21 42.30 44.75
100 37.17 37.45 36.19 42.64 44.55
Table 3: Impact of the vocabulary size on the downstream accuracy. The accuracy scores are averaged over the libraries and tokenizer algorithms
To provide a holistic performance overview, we assess the impact of the tokenizers on the model performance based on various tasks and analyze the results from different perspectives instead of providing solely a single aggregated metric.
This holistic evaluation is presented in Fig. 4, which expresses how often a tokenizer is among the top-1/-2/-3 performing tokenizers, Fig. 7, which shows the distribution of the accuracy scores across tasks, and in Table 9 which summarizes the average performances across all tasks. We dissect the results from different perspectives in the following subsections.
Monolingual Tokenizer
Among the monolingual tokenizers, there can be significant performance differences. For instance, Fig. 3(a) depicts that the GPT-2 tokenizer is overall most often across the top-$k$ (with $k\in\{1,2,3\}$) performing tokenizers, whereas the EN-BPE-HF-32 tokenizer is never the best-performing tokenizer. Though the count-based metric already provides a first overview of the model performance, it does not illustrate the difference in performance.
Therefore, we also looked at the accuracy distribution across all tasks that employ accuracy as the evaluation metric. Fig. 6(a) demonstrates that the EN-BPE-SP-32 tokenizer, on average, is across all tasks among the best-performing tokenizers and has the highest median value. At the same time, it also expresses a large variance across tasks compared to other tokenizers such as EN-BPE-HF-32, implying that for a set of tasks, the tokenizer performs very well, while on other tasks, it performs comparably poorly.
Aggregated metrics provide a good impression of the overall performance but do not demonstrate potentially large performance differences for single tasks. Therefore, we listed in Table 11 the performance difference across all tasks between the best-performing tokenizer and worst-performing tokenizer on this task. The performance difference is huge for various tasks, which is especially interesting since only the tokenizer has been changed while the model configuration is the same.
Overall, the GPT-2 tokenizer and the EN-BPE-SP-32 tokenizer are strong tokenizers for monolingual English models. Interestingly, SentencePiece’s implementation of BPE with a vocabulary size of 32k has been used for LLaMA2. Our results provide empirical and transparent evidence that both tokenizers/tokenizer settings are reasonable choices for training monolingual English models.
Multilingual Tokenizer
Our multilingual tokenizer experiments further support that different tokenizers can express large performance differences. Based on our count-based metric (Figure 3(b)), the differences across tokenizers are even larger than for monolingual tokenizers. While Multi-UNI-SP-100 and Multi-UNI-SP-82 are most often among the strongest tokenizers, Multi-BPE-HF-32 is only a single time across the top-performing tokenizers.
The accuracy distributions depicted in Figure 6(b) further show that both the tokenizers mentioned above are among the best-performing tokenizers and that specific tokenizers such as the already mentioned tokenizer perform worse. Figure 6(b) additionally demonstrates that the GPT-2 tokenizer performs poorly. Compared to monolingual models, we can observe that all tokenizers express a similar variance across tasks.
In reference to Table  11, the analysis of single tasks reveals that for multilingual tokenizers, the performance difference between tasks can be huge.
Multi-UNI-SP-100 and Multi-UNI-SP-82 are excellent choices, while employing a pre-trained GPT-2 tokenizer should be omitted to train multilingual models.
5.2.1 Impact of the Tokenizer Library
Table 2 demonstrates that BPE-SP, on average, outperforms BPE-HF across all tasks. In the multilingual setting, BPE-SP outperforms BPE-HF for all languages except for Spanish. The performance differences might be attributed to the differences in implementation details of the tokenisers’ pre-and postprocessing, which could affect the vocabulary creation (see Section 4.2) and, consequently, the downstream performance.
5.2.2 Impact of the Tokenizer Algorithm
Table 2 shows that for English tokenizers and models, BPE works better compared to Unigram. For multilingual models and tokenizers, we can see that for Spanish, German and French unigram can result in an improvement in performance. German and the romance languages are generally inflected languages, while English, although primarily analytic, does have some agglutinative features. Recent work showed Unigram tokenizers are better able to find morphologically accurate segmentations Bostrom and Durrett (2020); Park et al. (2021) which might help these languages.
5.2.3 Impact of the Tokenizer Vocabulary
Analyzing the impact of the vocabulary size revealed that in the monolingual English setting, the smaller/medium-sized vocabulary performs better (Table 3) whereas in the multilingual setting, larger vocabulary sizes result in better downstream performance(Table 3). This is in line with the vocabulary size of current English LLMs such as LLama.
5.3 Computational Costs
Even with an identical vocabulary size, two trained tokenizers most likely split a text passage in sequences of different lengths, resulting in different computation costs when processed by a model.
Our analysis revealed the following two insights. Firstly, increasing the vocabulary size decreases computational costs (Fig. 3). This implies that the reduction in the overall number of tokens represented by lower fertility (Fig. 1) compensates for the additional computational costs that a larger vocabulary size introduces in the model. In contrast, during inference, the computational costs increase with increasing vocabulary size and at least a vocabulary size of 50K, i.e., solely the forward pass without the backward pass does not compensate for the additional computational costs introduced by a larger vocabulary size in the model.
Secondly, the computational training costs for multilingual documents are significantly lower for multilingual tokenizers than monolingual English tokenizers for training (Fig. 2(a)) and for inference (in Appendix Fig. 8). In extreme cases, up to 68% additional training costs (see EN-UNI-SP-100352 vs. Multi-UNI-SP-50176 in Appendix Table 10 and Fig. 2(c)) as it is the case for training. At the same time, the difference in computational costs between multilingual and monolingual English tokenizers for processing English documents is much lower (Fig. 2(b)). A complete overview of the computational costs can be found in Table 10.
6 Correlation Between Intrinsic And Extrinsic Tokenizer Performance
Figure 5: Correlation of fertility/parity scores and downstream task performance for all five languages. We evaluated English models on English tasks, whereas multilingual models are evaluated across all four non-English tasks.
This section investigates a possible predictive relationship of intrinsic tokenizer metrics (fertilty and parity) to the extrinsic model downstream performance.
As highlighted in the correlation heatmaps in Fig. 5, we find that there is no distinct correlation across all tasks and languages, demanding a more granular analysis. While for non-English tasks, we mainly observe a correlation between low fertility and higher downstream performance, the non-English tasks yield seemingly random positive and negative correlations.
Taking the fertility trends with varying vocabulary sizes (see Fig. 1) into consideration, we hypothesize that fertility only correlates with downstream performance in certain language-specific vocabulary size limits. For the English language, the tokenizers already provide low, close-to-convergence fertility scores for vocabulary sizes of 32k tokens. While additional tokens yield only minute fertility improvements, we presume that they do not capture morphological segmentations and, thus, can harm downstream performance and significantly increase the computation costs (see Section 5.3) in the end.
In contrast, for multilingual tokenizers, we observe significant fertility improvements with increasing vocabulary sizes. Due to the larger thesaurus induced by the additional languages, the tokenizer requires a larger vocabulary to allow a model to perform convincingly on all languages. Therefore, only within the non-convergence vocabulary range, we achieve a strong, negative correlation between fertility and downstream performance with varying vocabulary sizes.
In conclusion, intrinsic tokenizer metrics such as fertility and parity need to be taken with a grain of salt and supposedly are only predictive of downstream model performance in certain bounds. Furthermore, the analysis showed that low fertility scores can be regarded as a necessary criterion but not as a sufficient one.
7 Conclusion
This work represents a fundamental step to a better understanding of the impact of the tokenizer on the models’ downstream performance. We have shown that training tokenizers with a balanced share across languages achieve comparable low fertility and parity scores across all languages, which has important implications. Lower fertility results in less computational costs and enables the model to learn long-range dependencies in limited context windows.
Furthermore, we highlight that the tokenizer choice can significantly impact the model’s downstream performance. We could show that it is beneficial to employ Unigram tokenizers with a large vocabulary size to train multilingual models. This is contrary to the current monolingual setting where BPE-based tokenizers are used together with a smaller/medium-sized vocabulary size. This also implies that insights obtained for monolingual settings are not transferable to multilingual settings and, therefore, require dedicated studies.
Finally, we could show that there is no clear correlation between intrinsic and extrinsic tokenizer performance, but the correlation is rather task-specific.
Acknowledgements
This work was funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D) as well as by the Federal Ministry of Education and Research of Germany and the state of North-Rhine Westphalia as part of the Lamarr-Institute for Machine, LAMARR22B. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC) as well as the Center for Information Services and High Performance Computing [Zentrum für Informationsdienste und Hochleistungsrechnen (ZIH)] at TU Dresden for providing its facilities for high throughput calculations.
Appendix A Corpora
Name Language #Words
Oscar DE 11.200.000.000
Oscar ES 11.200.000.000
Oscar EN 11.200.000.000
Oscar IT 11.200.000.000
Oscar FR 11.200.000.000
Pile DE 13.838.432
Pile ES 21.990.512
Pile EN 4.334.313.669
Pile IT 7.946.402
Pile FR 15.857.811
RedPajama DE 143.907.461
RedPajama ES 112.950.000
RedPajama EN 4.663.646.781
RedPajama IT 137.802.711
RedPajama FR 139.749.147
RedPajama Code 2.052.228.788
Misc DE 600.844.912
Misc ES 186.934.269
Misc EN 1.337.030.904
Misc IT 19.810.753
Misc FR 211.147.445
Total 70.000.000.000
Table 4: Overview of the multilingual 70B words dataset with language, number of sampled words
Name Language # words
Oscar EN 56.000.000.000
Pile EN 4.893.724.288
RedPajama EN 5.308.974.750
RedPajama Code 2.299.301.635
Misc EN 1.497.999.327
Total 70.000.000.000
Table 5: Overview of the English 70B words dataset with language, number of sampled words
Our web documents in the corpora consist of Oscars [^1] https://oscar-project.org/ Abadji et al. (2021), that were generated by the ungoliant pipeline [^2] https://github.com/oscar-project/ungoliant based on three Common Crawl WET Archives (2022-27, 2022-49 and 2023-14).
The curated datasets consist of The Pile Gao et al. (2020a), RedPajama Computer (2023), and single datasets that do not belong to a collection. From the Pile subcorpora, we selected: Phil Archive, PMC Abstracts, PMC Extracts, OpenWebText, NIH Exporterm, and Free Law Opinions V2. From RedPajama we use: ArXiv, Books, Github, StackExchange, and Wikipedia.
The remaining datasets are:
Hyper-Parameter Value(s)
model_type Unigram |||| BPE
vocab_size 33k ||||50k |||| 82k |||| 100k
character_coverage 0.9999
split_by_number True
allow_whitespace_only_pieces True
add_dummy_prefix True
user_defined_symbols_1 <s>, </s>,<pad>,<eod>
user_defined_symbols_2 …, <placeholder_255>
byte_fallback True
max_sentence_length 4192
normalization_rule_name NFKC
train_extremely_large_corpus True
remove_extra_whitespaces False
split_by_whitespace True
Table 6: Overview of the SentencePiece options that we used for the training of our tokenizers.
Hyper-Parameter Value(s)
model_type BPE
vocab_size 33k ||||50k |||| 82k |||| 100k
limit_alphabet 512
use_nfkc_normalizer True
use_lowercase_normalizer False
use_strip_accents_normalizer True
pre_tokenizer ByteLevel, Digits
Table 7: Overview of the Sentencepiece Options that we used for the training of our tokenizers.
  1. 1.
    All the News V2.0333https://metatext.io/datasets/all-the-news-2.0 is a corpus of newspaper articles crawled from over 26 different publications from January 2016 to April 1, 2020.
  2. 2.
    Bundestag - Plenarprotokolle [^4] https://www.bundestag.de/dokumente/protokolle/plenarprotokolle comprises transcripts of sessions of the German Bundestag.
  3. 3.
    Bundesgerichtshof - Entscheidungen [^5] https://www.bundesgerichtshof.de/DE/Entscheidungen/entscheidungen_node.html is a collection of decisions of the German Federal Court.
  4. 4.
    CoStEP [^6] https://pub.cl.uzh.ch/wiki/public/costep/start is a cleaned-up and corrected version of the EuroParl corpusGraën et al. (2014). Koehn (2005)
  5. 5.
    DCEP [^7] https://joint-research-centre.ec.europa.eu/language-technology-resources/dcep-digital-corpus-european-parliament_en is a companion corpus to CoStEP, containing documents published by the European Parliament. Hajlaoui et al. (2014)
  6. 6.
    DNB Dissertations [^8] https://www.dnb.de/DE/Professionell/Services/Dissonline/dissonline_node.html is a collection of dissertations from the Deutsche Nationalbibliothek.
  7. 7.
    MAREC/IREC [^9] https://researchdata.tuwien.ac.at/records/2zx6e-5pr64: The MAtrixware REsearch Collection / The Information retrieval facility Research Collection is a patent corpus of over 19 million documents from the EP, WO, US, and JP patent offices.
  8. 8.
    Medi-Notice [^10] https://pub.cl.uzh.ch/wiki/public/pacoco/medi-notice is part of the Zurich Parallel Corpus Collection. It is a multilingual corpus compiled from information leaflets for medications and pharmaceutical products published by the Swiss Agency for Therapeutic Products.Graën et al. (2019)
  9. 9.
    Swiss Policy [^11] https://pub.cl.uzh.ch/wiki/public/pacoco/swiss_legislation_corpus contains documents of the Swiss Legislation Corpus Höfler and Piotrowski (2011)
  10. 10.
    OpenSubtitles 2018121212https://opus.nlpl.eu/OpenSubtitles-v2018.php [^13] https://www.opensubtitles.org/de/index.cgi is a collection of translated movie subtitles. Lison and Tiedemann (2016)
Appendix B Tokenizer
In our experiments, we focused on the Huggingface tokenizer library Moi and Patry (2023) and the SentencePiece library Kudo and Richardson (2018). We use the standard settings of the SentencePiece library if not stated otherwise in Table 6. For the HuggingFace tokenizer library Table 7 shows where we deviated from the standard values.
Appendix C LLM Architecture and Hyperparameters
Regarding the training architecture of our 2.7B parameter models, we followed closely the architecture of GPT-3 Brown et al. (2020a). An overview of the used architecture details and hyperparameters is given in Table 8.
parameter value
# Hidden Dimension 2560
# Layers 32
# Attention-Heads 32
Sequence-Length 2048
Optimizer Adam
[…]
Learning rate 1.6e-4
Learning rate decay Cosine
Precision BF16
FlashAttention 2.0
Position-Embeddings Rotary
Table 8: Overview of the LLM hyperparameters that we used for the training.
Appendix D Intrinsic Tokenizer Evaluation
Besides studying the overlap of the same algorithm on the same thesaurus, we were also interested in vocabulary overlaps across algorithms and thesauruses see Fig. 6. What we can observe is that multilingual vocabulary and English vocabulary have a rather small overlap between 24% and 34% that remains similar across increasing vocabulary sizes. Across algorithms, we can see that Unigram and BPE of SentencePiece have a slightly higher overlap than Unigram of SentencePiece and BPE of Huggingface. We think this might be due to library-specific preprocessing steps and more similar hyperparameters.
Figure 6: Vocabulary overlap between the examined tokenizers
Appendix E Extrinsic Tokenizer Evaluation
Accuracy
Model EN MULTI
BPE-HF-100 45.60 38.50
BPE-HF-32 45.34 38.12
BPE-HF-50 45.85 38.15
BPE-HF-82 44.97 38.67
BPE-SP-100 46.06 38.92
BPE-SP-32 47.06 38.01
BPE-SP-50 46.85 38.41
BPE-SP-82 45.25 38.64
UNI-SP-100 45.97 39.34
UNI-SP-32 46.62 38.18
UNI-SP-50 45.81 38.66
UNI-SP-82 45.65 39.21
Table 9: Average accuracy of monolingual and multilingual tokenizers across all downstream tasks
(a) Monolingual tokenizers
(b) Multilingual tokenizers
Figure 7: Accuracy distribution for mono-/multilingual tokenizers across mono-/multilingual tasks. The "+" represents the average performance across all tasks.
(a) Non-English documents.
(b) English documents.
(c) German documents.
Figure 8: Average compute (GFLOPs) required to process a single word within (a) multilingual, (b) English, and (c) German documents within an inference pass.
E.1 Performance Gains on Task-Level
In table Table 11, we show how much influence the choice of a tokenizer has on each downstream task. Therefore, we compare the results of the best- and worst-performing models for each task in our experiments. Performance increasements reach from +5,3% up to 380,9%, again showing the importance of choosing the right tokenizer on the downstream performance.
Model Non-English English German
EN-BPE-HF-33 3,8 2,32 4,52
EN-BPE-HF-50 3,79 2,38 4,45
EN-BPE-HF-82 3,88 2,55 4,51
EN-BPE-HF-100 3,96 2,67 4,58
EN-BPE-SP-33 3,86 2,37 4,66
EN-BPE-SP-50 3,89 2,42 4,68
EN-BPE-SP-82 4,02 2,59 4,78
EN-BPE-SP-100 4,11 2,71 4,84
EN-UNI-SP-32 4,01 2,36 4,73
EN-UNI-SP-50 4,02 2,42 4,75
EN-UNI-SP-82 4,12 2,59 4,83
EN-UNI-SP-100 4,21 2,71 4,88
MULTI-BPE-HF-33 2,71 2,46 3,04
MULTI-BPE-HF-50 2,7 2,5 3,01
MULTI-BPE-HF-82 2,8 2,65 3,09
MULTI-BPE-HF-100 2,88 2,76 3,17
MULTI-BPE-SP-33 2,68 2,55 2,99
MULTI-BPE-SP-50 2,67 2,57 2,95
MULTI-BPE-SP-82 2,76 2,72 3,03
MULTI-BPE-SP-100 2,85 2,82 3,1
MULTI-UNI-SP-33 2,68 2,55 2,94
MULTI-UNI-SP-50 2,66 2,58 2,91
MULTI-UNI-SP-82 2,76 2,73 2,99
MULTI-UNI-SP-100 2,84 2,83 3,07
Table 10: Computational costs per word (GFLOPs) for different tokenizers.
Task Metric Worst Best Difference
EN arc challenge acc 0,21 0,27 +28,4 %
arc easy acc 0,50 0,59 +18,6 %
boolq acc 0,56 0,61 +8,9 %
cb acc 0,30 0,54 +76,5 %
cb f1 0,18 0,36 +94,9 %
copa acc 0,67 0,73 +9,0 %
hellaswag acc 0,34 0,41 +19,7 %
lambada openai acc 0,47 0,56 +20,0 %
mathqa acc 0,22 0,24 +8,8 %
mnli acc 0,35 0,37 +7,7 %
mnli mismatched acc 0,35 0,38 +9,4 %
mrpc acc 0,54 0,69 +26,1 %
mrpc f1 0,66 0,81 +22,7 %
gnad10 acc 0,15 0,43 +187,7 %
gnad10 f1 0,07 0,35 +380,9 %
gnad10 recall 0,14 0,42 +207,6 %
pawsx es acc 0,48 0,56 +17,2 %
piqa acc 0,67 0,72 +7,2 %
pubmedqa acc 0,45 0,59 +32,1 %
race acc 0,31 0,34 +11,8 %
rte acc 0,53 0,57 +8,9 %
sst acc 0,51 0,75 +48,2 %
triviaqa acc 0,02 0,05 +171,3 %
wic acc 0,50 0,55 +11,4 %
winogrande acc 0,49 0,56 +12,4 %
wnli acc 0,42 0,54 +26,7 %
wsc acc 0,36 0,61 +70,3 %
MULTI xstory cloze es acc 0,49 0,60 +24,3 %
mgsm de acc 0,00 0,03 -
lambada de acc 0,11 0,33 +186,4 %
lambada es acc 0,17 0,32 +91,8 %
lambada fr acc 0,20 0,38 +90,5 %
lambada it acc 0,16 0,37 +132,3 %
wino de acc 0,50 0,61 +20,9 %
xcodah de acc 0,27 0,42 +52,1 %
xcodah es acc 0,28 0,43 +52,1 %
xcodah fr acc 0,27 0,39 +43,7 %
xcsqa de acc 0,19 0,30 +56,4 %
xcsqa es acc 0,19 0,29 +57,3 %
xcsqa fr acc 0,17 0,29 +69,2 %
xnli de acc 0,32 0,48 +51,0 %
xnli en acc 0,49 0,52 +5,3 %
xnli es acc 0,34 0,48 +41,0 %
xnli fr acc 0,37 0,49 +30,3 %
xstance de acc 0,49 0,52 +6,0 %
xstance de f1 0,34 0,47 +41,2 %
xstance de recall 0,49 0,52 +7,7 %
Table 11: Differences between our best and worst performing tokenizers per task.